In triangle ABC prove that cot A/2+cot B/2+cot c/2=a+b+c/b+c-a cot (B/2)= a+b+c/ a+b-c cot(C/2)

Grade:11

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please write the Question in a standard form or attach an image of the question

Regards
Vikas TU
4 years ago
Dear student 
Dear student 
A/2 + B/2 = pi/2 - C/2 
Cot(A/2+ B/2) = Cot(pi/2-C/2 ) 
= tanC/2 = 1/Cot(c/2)
Apply cot(x+y) = (Cotxcoty-1)/(cotx+coty)
 
cot A/2+cot B/2+cot c/2=a+b+c/b+c-a cot (B/2)= a+b+c/ a+b-c cot(C/2)
Good Luck